OpenStudy (anonymous):

Solve for the variable. 4.1b = -20.5

OpenStudy (anonymous):

Isolate b by dividing both sides of your equation by 4.1.

OpenStudy (anonymous):

I got -16.4 which is wrong...

OpenStudy (anonymous):

\[\frac{ 4.1b }{ 4.1 } = \frac{ -20.5 }{ 4.1 }\]

OpenStudy (anonymous):

\[b = \frac{ -20.5 }{4.1} \]

OpenStudy (anonymous):

You are subtracting, you need to divide.